An investor invests $4000 to buy 200 shares of Sand Corporation, which has an expected return of 12%; $5000 to buy 300 shares of Water Corporation, with an expected return of 13%; and $6000 to buy 400 shares in Beach Corporation, with an expected return of 15%. What is the expected return on this portfolio? O 13.3% 13.0% O 13.5% 13.3%
Answer #1
Expected return on the given portfolio=
Sum of (Proportional $ Weights of individual investment* Their Expected returns)
ie. (Wt.Sand Corpn.shares*ER(SC))*(Wt.Water Corpn.shares*ER(WC))*(Wt. Beach Corpn.shares*ER(BC))
Total $ investment= 4000+5000+6000=
15000
Wt. of Sand Corporation shares in the total portfolio=4000/15000= 4/15
Wt. of Water Corporation shares in the total portfolio=5000/15000= 5/15
Wt. of Beach Corporation shares in the total portfolio=6000/15000= 6/15
Now,
the Expected return on the given portfolio=(4/15*12%)+(5/15*13%)+(6/15*15%)=
13.53%
So, the ANSWER is
C. 13.5%